HIGH RESOLUTION EMISSION SPECTROSCOPY OF THE VIBRATION-ROTATION BANDS OF HBO AND HBS.

G. LI^{a}, \underline{R.S. RAM}, R.J. HARGREAVES, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; P.F. BERNATH, Department of Chemistry and Biochemistry, Old Dominion University, Norfolk, VA 23529 USA; Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; and H. LI, State Key Lab of Theoretical and Computational Chemistry, Jilin University, Changchun City, China, 130023.

The vibration-rotation spectra of HBO and HBS have been investigated at high resolution using a Fourier transform spectrometer. The HBO molecules were produced in a high temperature furnace from the reaction of H_{2}O vapor with boron by heating a mixture of crystalline boron and boron oxide (B_{2}O_{3}) at a temperature ∼1350^{∘}C. The spectra were recorded in the 1100–2200 cm^{−1} and 1700–4000 cm^{−1} wavenumber regions covering the ν_{3} and ν_{1} fundamentals, respectively. In total 24 vibrational bands involving 30 vibrational levels of H^{11}BO and 12 bands involving 18 levels of H^{10}BO have been rotationally analyzed. After combining the existing microwave and infrared measurements, the absolute term values have been determined for a number of vibrationally-excited states of H^{11}BO and H^{10}BO.

The HBS molecules were formed by the reaction of CS_{2} and water vapor with crystalline boron at a temperature ∼1300^{∘}C. The spectra were recorded in the 850–1500 cm^{−1} and 1750–4000 cm^{−1} wavenumber regions covering the ν_{3} and ν_{1} frequency regions. In total 29 vibrational bands involving 33 vibrationally-excited levels of H^{11}BS and 9 bands involving 12 vibrational levels of H^{10}BS have been analyzed. The fitted spectroscopic parameters agree very well with the results of our ab initio calculations. L-resonance interactions observed between the 02^{0}0 (Σ) and 02^{2}0 (Δ) levels of HBO and HBS were analyzed using a 2×2 matrix to yield deperturbed constants.

aCurrent address: Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, MS#50, 60 Garden St., Cambridge, MA, 02138, USA.